# Significance of soft-scale breaking on primordial black hole production in Coleman-Weinberg type supercooling-phase transition

### v1

@@ I Introduction
CSI extensions of the Standard Model (SM), which encode the CW mechanism, are attractive since they not only provide a potential resolution to the hierarchy problem in particle physics via the dimensional transmutation, but also have been successfully applied to cosmological contexts, including the so-called CW-type small-field inflation [14, 15] and the generation of a stochastic gravitational wave (GW) background [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], hence making them both theoretically and cosmologically well-motivated.
Recently, CSI theories have received considerable attention in a view of the PBH formation via the mechanism of delayed vacuum transitions [28, 29, 30, 31, 32, 33, 34]. According to semi-analytical estimates [28], in the ultra-supercooling regime ($\alpha\gtrsim 10^{2}$), explaining the observable dark matter via PBHs requires the parameter $\beta/H$ to lie within the range of $\sim 6.4-6.9$. If $\beta/H$ falls below this range, PBHs would be overproduced and rapidly dominate the energy density of the universe. Conversely, if $\beta/H$ exceeds this range, corresponding to a rapid phase transition, the resulting PBH abundance would be insufficient to account for dark matter.
@@ III.2 The case with soft-scale breaking mass
We numerically evaluate $S_{3}(T)/T$ and $\beta_{H}$ as a function of $g_{D}$, $v_{\phi}$, and $m_{0}$. See Fig. 1, for some reference points in the parameter space. The plots in the figure have been generated without using the approximated formulae in Eqs. (19), (25), and (27). The figure indicates the numerical validity of the semi-analytical expressions given in them.
Figure 1: Left: Plots of the bounce action $S_{3}(T)/T$ in Eq. (25) with a negative (denoted as “Neg.”) and positive (denoted as “Pos.”) soft-scale breaking terms, as a function of $T$, for reference values of $g_{D}$ and $v_{\phi}$. Right: Variations of $\beta_{H}$ with respect to the coupling $g_{D}$ for different mass choices and a fixed $v_{\phi}$ consistently with the PBH-dark matter interpretation (See Eq. (44)).

### v2

@@ I Introduction
CSI extensions of the Standard Model (SM), which encode the CW mechanism, are attractive since they not only provide a potential resolution to the hierarchy problem in particle physics via the dimensional transmutation, but also have been successfully applied to cosmological contexts, including the so-called CW-type small-field inflation [14, 15] and the generation of a stochastic gravitational wave (GW) background [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], hence making them both theoretically and cosmologically well-motivated.
Recently, CSI theories have received considerable attention in a view of the PBH formation via the mechanism of delayed vacuum transitions [30, 31, 32, 33, 34, 35, 36, 37, 38]. According to semi-analytical estimates [30], in the ultra-supercooling regime ($\alpha\gtrsim 10^{2}$), explaining the observable dark matter via PBHs requires the parameter $\beta/H$ to lie within the range of $\sim 6.4-6.9$. If $\beta/H$ falls below this range, PBHs would be overproduced and rapidly dominate the energy density of the universe. Conversely, if $\beta/H$ exceeds this range, corresponding to a rapid phase transition, the resulting PBH abundance would be insufficient to account for dark matter.
@@ III.2 The case with soft-scale breaking mass
We numerically evaluate $S_{3}(T)/T$ and $\beta_{H}$ as a function of $g_{D}$, $v_{\phi}$, and $m_{0}$. See Fig. 1, for some reference points in the parameter space. It is worth noting that the approximate formulae in Eqs. (19), (25), and (27), are introduced merely to provide an intuitive understanding of the impact of the soft-scale breaking terms. The plot in this figure, as well as the subsequent results, are obtained from full numerical computations rather than relying on these approximations.
Figure 1: Left: Plots of the bounce action $S_{3}(T)/T$ with a negative (denoted as “Neg.”) and positive (denoted as “Pos.”) soft-scale breaking terms, as a function of $T$, for reference values of $g_{D}$ and $v_{\phi}$. Right: Variations of $\beta_{H}$ with respect to the coupling $g_{D}$ for different mass choices and a fixed $v_{\phi}$ consistently with the PBH-dark matter interpretation (See Eq. (44)).
@@ VII Summary and discussion
The three-dimensional effective theory analysis [106, 107, 38, 108] may yield quantitatively different results. A thorough reanalysis incorporating these effects, as well as recent refinements of PBH abundance estimates accounting for trapped false vacuum domains, nucleation time fluctuations of multiple bubbles, horizon-scale fluctuations during thermal inflation, and higher-order corrections to the bubble nucleation rate [109, 110, 111], lies beyond the scope of this work and will be pursued in future studies. Here, we focus on the impact of soft-scale breaking on PBH formation using a simpler approach, which we expect to capture the essential physics.